OpenStudy (anonymous):

200-gallon tank that is full of water needs to be drained and it drains at a constant rate of 15-gallons per minute. Write a linear function representing the number of gallons of water left in the tank after x minutes. About how long will it take to empty the tank

OpenStudy (anonymous):

g = -15m + 200 15m = 200 =>m = 200/15 = 13.3 min